OpenStudy (anonymous):

57 rainbowink for each problem do the following. a define a verialble b wright an equation to modle the situation. 1. the math club made a total of $1,155 from selling 220 t-shirts. how much did they sell each t shirt for. 2. a youth basketball league recently compleated regestration.the league was able to form 15 teams of 11 players each how many players registered? 3.at the local carnaval you bought 15 ride tickets and a sovinier stuffed animal for $12.00 you spent a total of $30.00 and 75 cents. how much did u spend on each ride tickets? 4. a local garden center sells 8 in

OpenStudy (anonymous):

sue will answer them couse shes boss.

OpenStudy (anonymous):

lol whats that supposed to mean? 1.) Divide 1155/220 = $5.25. 2.) Multiply 15 x 11 = 165 players 3.) I would first subtract the stuffed animal from the total money she spent. 30.75 - 12.00 = 18.75 Then I would divide. 18.75/15 = $1.25 per ticket

OpenStudy (anonymous):

I don't understand the last question. :)

OpenStudy (anonymous):

dats okay (: somone else answer it?

OpenStudy (anonymous):

I don't think its complete thats what I mean ;)